OpenStudy (anonymous):

lim Cot(2x) Sin(6x) (l'hospital's rule section) x -> 0

OpenStudy (dumbcow):

to apply L'hopitals rule you need to put it in fraction form \[\rightarrow \frac{\cos(2x) \sin(6x)}{\sin(2x)}\] now differentiate, use product rule for numerator (fg)' = f'g + fg'

OpenStudy (anonymous):

what he said

OpenStudy (anonymous):

take the derivative of the top.. and the bottom separately [-2sin(2x)sin(6x) + 6cos(2x)cos(6x)] / -2cos(2x)

OpenStudy (anonymous):

= [-sin(2x)sin(6x) + 3cos(2x)cos(6x)] / -cos(2x)

OpenStudy (anonymous):

= 0 + 3 * 1 / - 1

OpenStudy (anonymous):

= -3

OpenStudy (anonymous):

actually answer is positive 3

OpenStudy (anonymous):

but i figured it all out now, ty both

OpenStudy (anonymous):

er yeah sorry should have been -6cos(6x)... equating to 3 srry

OpenStudy (anonymous):

no, the deriv of sin is cos... not -cos. problem was in the denom. but moot point.. i got it!!

OpenStudy (anonymous):

i will give you medal anyway :D

OpenStudy (anonymous):

errrrr sorry late night i meant to say 2cos(2x) :) :) read the wrong thing

OpenStudy (anonymous):

glad this isn't a test lol thank ye good luck :0

OpenStudy (anonymous):

:)